1. The area of a rectangular room is 300 ft2. The perimeter is 70 ft. What are the dimensions of the room?
i tried to do this, but it doesnt work!!!!
100x3, 2x150, 1x300, 5x60, 6x50
thx!

Answer by ankor@dixie-net.com(22740)

1. The area of a rectangular room is 300 ft2.
The perimeter is 70 ft. What are the dimensions of the room?
:
The perimeter:
2L + 2W = 70
Simplify, divide by 2
L + W = 35
W = (35-L); use this form for substitution
:
The area:
L * W = 300
replace W with (35-L)
L*(35-L) = 300
-L^2 + 35L = 300
-L^2 + 35L - 300 = 0
Easier to factor if L^2 is positive, multiply by -1
L^2 - 35L + 300 = 0
Factor this to
(L-15)(L-20) = 0
Two solutions
L = 15
L = 20 ft
choose the larger value for L, 20 ft, find W
W = 35 - 20
W = 15 ft is the width
:
:
See if that is true
2(20) + 2(15) = 70